# HOI-Swap: Swapping Objects in Videos with Hand-Object Interaction Awareness

### v1

@@ 1 Introduction
Consider a video depicting the process of a human hand picking up a kettle and moving it around (Figure 1). What if we want to replace the hand-interacting object in this scene with another item specified by the user—perhaps a differently-shaped kettle, a bottle, or a bowl? This capability—to swap the in-contact object in a video with another, while aligning with the original video’s content—is crucial for enhancing various real-world applications. Such functionality can transform entertainment, allowing users to create novel video content without the need to re-record or engage in labor-intensive manual editing. Additionally, it holds significant promise in robotics, where recent results suggest generative models can reduce the reliance on manually collected task-specific visual data and thereby enable large-scale robot learning kapelyukh2023dall ; yang2023learning .
@@
Second, consider the challenge of spatial alignment with the original video. The reference object might appear in any arbitrary pose; for instance, in Figure 2 (b), the reference kettle handle is on the left in the reference image, but for realistic interaction, the generated content needs to reposition the kettle handle to the right, where the hand is poised to grasp it. However, current approaches do not offer this level of control, as evidenced by the results from a hand insertion approach Affordance Diffusion (Afford Diff) ye2023affordance in Figure 2 (b). Despite being adept at generating a hand interacting with the given kettle, it does not ensure correct object placement to align with the hand and scene context in the original image, lacking spatial alignment capability.
@@ 3.1 Task Formulation and Framework Overview
We create a fully self-supervised training approach, necessitated by the impracticality of collecting paired videos $(\mathcal{V},\mathcal{V}^{\star})$. We pose the problem as a video inpainting task. Specifically, from the original training video $\mathcal{V}$ and mask $\mathcal{M}$, we derive a masked video $\mathcal{V}^{m}$, accompanied by a set of reference object images $\{\mathcal{I}^{ref}\}$. During training, the model takes the masked video sequence $\mathcal{V}^{m}$ and one randomly chosen object image $\mathcal{I}^{ref}$ that originates from elsewhere in the training video to reconstruct $\mathcal{V}$. During inference, the model is given a bounding box-masked sequence combined with various object images to test its swapping capability.
As outlined in Section 1, the task presents three main challenges: HOI awareness, spatial and temporal alignment. In response, we propose a two-stage framework that decomposes the complexity. The first stage integrates the reference object into a single frame, targeting HOI awareness and establishing spatial alignment initially. The second stage propagates the one-frame edit across the entire video sequence, focusing on temporal alignment and carrying forward the spatial alignment from the first stage. This structured approach effectively lightens the model’s generation load. In terms of which frame to select from the input video for stage-I edits, we automate the process by adopting an off-the-shelf hand-object detector shan2020understanding to identify frames with hand-object contact, from which we randomly select one for editing. See Supp. C.3 for a detailed analysis. For greater applicability and flexibility, each stage is trained separately in a self-supervised manner, using the original object image as $\mathcal{I}^{ref}$. Consequently, the stage-I model not only plays a crucial role in the video editing pipeline but also serves as a standalone image editing tool. The full pipeline is illustrated in Figure 3.
@@ 3.2 Stage I: HOI-aware Object Swapping in One Frame
  • Object augmentation: Directly forming a training pair ($\mathcal{I}_{i}^{obj}$, $\mathcal{I}_{i}^{m}\rightarrow\mathcal{I}_{i}$) is suboptimal as it does not reflect the variability expected in real-world applications. At test time, the reference object may appear in any pose, orientation, or size. To bridge this gap between training and testing scenarios, we apply strong augmentation techniques: (1) Spatially, we enhance the object’s diversity by applying random flips, rotations, and perspective transformations; (2) Temporally, instead of using the object image from its original frame $i$, we introduce variability by randomly selecting an alternate frame, $\mathcal{I}_{i^{\prime}}^{obj}$, from all available frames in the source video that contain the object. Finally, for preservation of the reference object’s identity and structural details, we collage the augmented reference object image $\mathcal{I}_{i^{\prime}}^{obj}$ onto the masked region in $\mathcal{I}_{i}^{m}$, producing $\mathcal{I}_{i}^{m^{\prime}}$.
@@ 3.3 Stage II: Controllable Motion-guided Video Generation
In stage II, given the first-stage edit $\mathcal{I}_{i}^{\star}$, source video $\mathcal{V}$, and its binary mask sequence $\mathcal{M}$, the objective is to generate a new video $\mathcal{V}^{\star}$ that propagates the single-frame edit across the remaining frames. For this purpose, we perform warping to transfer pixel points from the edited frame to the remaining frames, resulting a sparse and incomplete video sequence. We then train a video diffusion network that learns to correct and fill in the gaps, completing the video editing process.
@@
  • Conditioning frame selection. To avoid limiting the model to generating videos based on a specific reference frame, such as the first frame as often used in existing image-to-video generative models, we randomly select a frame $\mathcal{I}_{c}$ from $\mathcal{V}$ as the conditioning frame. This any-frame-conditioning mechanism brings additional flexibility during inference (for detailed discussion, see Supp. C.3).
@@
We vary the sparsity level of sampled motion points, $r$ from 0 to 100 during training, preparing the model for a range of scenarios—from no motion guidance to full motion information. During inference, we demonstrate that the model is capable of generating video sequences with varying levels of motion alignment to the original video (see Section 4.2 and Supp. C.3 for further discussion). Thus, our stage-II training goal is to take the warped yet incomplete video sequence ($\mathcal{V}^{warp}$) as input, then learn to fill in the gaps and smooth over discontinuities to reconstruct the full video ($\mathcal{V}$).
@@
Our video LDM involves two conditional signals in addition to the standard $\mathbf{z}_{t}$ (a noised version of $\mathbf{z}$) and diffusion time step $t$: (1) The warped video sequence $\mathcal{V}^{warp}\in\mathbb{R}^{N\times 3\times H\times W}$, is processed similarly as $\mathcal{V}$, encoded by $\mathcal{E}$, producing a latent feature $\mathbf{z}^{w}\in\mathbb{R}^{N\times 4\times H/8\times W/8}$. $\mathbf{z}^{w}$ is then concatenated channel-wise with $\mathbf{z}$ before being fed into the denoising network $\epsilon_{\theta_{2}}$. (2) The selected frame $\mathcal{I}_{c}$, besides serving as the anchor in the warping process, is also encoded by CLIP radford2021learning , yielding $d^{c}\in\mathbb{R}^{768}$ that provides high-level contextual guidance and is incorporated in $\epsilon_{\theta_{2}}$ via cross-attention.
@@ 4 Experiments
To comprehensively evaluate HOI-Swap, we consider both the image editing (accomplished by the stage-I model) and video editing (facilitated by the entire pipeline) task. We describe experimental setup in Section 4.1 and present editing results along with an ablation analysis in Section 4.2.
@@ 4.1 Experimental Setup
Datasets Our training leverages two large-scale egocentric datasets, HOI4D liu2022hoi4d and Ego-Exo4D grauman2023ego , which feature abundant HOIs, making them particularly suitable for exploring this problem. In total, the stage-I model is trained on 106.7K frames and the stage-II model uses 26.8K 2-second video clips for training. The evaluation set for image editing includes 1,250 source images, each paired with four reference object images, resulting in a total of 5,000 edited images. The evaluation set for video editing is composed of 25 source videos, also combined with four object images each, yielding 100 unique edited videos. Note that: (1) For image editing evaluation, we include both hand-present scenarios, which challenge the model’s HOI-aware capabilities, and non-hand-present scenarios to evaluate its general object-swapping capability. (2) For video editing evaluation, in addition to HOI4D videos, we include challenging in-the-wild videos from EPIC-Kitchens damen2018scaling and third-person videos from TCN Pouring sermanet2017time —datasets not used in training—to assess the model’s zero-shot generalization across diverse scenarios. See Supp. B.1 for complete dataset descriptions.
Baselines For image editing, our stage-I model is compared with two strong image inpainting approaches, PBE yang2023paint and AnyDoor chen2023anydoor , along with a hand-insertion approach Afford Diff ye2023affordance in the HOI domain. For video editing, we consider the following baselines: (1) applying the best image editing baseline to each frame of the video; (2) integrating the best image editing baseline with AnyV2V ku2024anyv2v , a recent video editing approach that takes one edited frame from any black-box image editing model as conditional guidance; (3) VideoSwap gu2023videoswap , the state-of-the-art approach in customized object swapping. Besides these object-image-guided approaches, we include a focused comparison with text-guided approaches for a thorough evaluation. See Supp. C.5 for the discussion.
Figure 4: Qualitative Results of HOI-Swap. We compare HOI-Swap with image (left) and video editing approaches (right). The reference object image is shown in the upper left corner of the source image. For image editing, HOI-Swap demonstrates the ability to seamlessly swap in-contact objects with HOI awareness, even in cluttered scenes. For video editing, HOI-Swap effectively propagates the one-frame edit across the entire video sequence while accurately following the source video’s motion, achieving the highest overall quality among all methods. We highly encourage readers to check Supp. C.1 and Supp. video for more comparisons.
@@ 4.2 Editing Results
Quantitative evaluation Table 1 compares the performance of HOI-Swap on both image and video editing tasks. For image editing, the two inpainting approaches, PBE yang2023paint and AnyDoor chen2023anydoor struggle to generate the hand correctly, as indicated by hand fidelity scores below 90%. Afford Diff ye2023affordance is developed for the HOI context and also benefits from HOI-specific training data in HOI4D. It produces high-fidelity hands but does not address the editing task adequately, resulting in low contact and hand agreement with the source image. Our proposed HOI-Swap excels in both aspects, achieving superior performance in agreement and fidelity scores. In the human evaluation, HOI-Swap is consistently favored over the other three approaches, achieving a preference rate of 72.1%. For video editing, HOI-Swap outperforms all competing methods across all metrics, including general video quality metrics (subject consistency and motion smoothness) and HOI edit metrics (contact agreement, hand agreement and hand fidelity). It significantly surpasses the three baseline approaches, winning the user study with a selection rate of 86.4%. Recognizing the inherent challenges of the video editing task, we included a survey question asking users if “All edits are of very low quality; none successfully swap the object.” This response was selected at an average rate of 10.9%. See Supp. C.2 for details.
@@
For video editing (Figure 4 right), the source video depicts a hand rotating a toy car. The per-frame approach struggles with temporal consistency of the object identity and produces unnatural HOI motions. AnyV2V ku2024anyv2v fails to propagate the first-frame edit through subsequent frames, leading to inconsistencies. VideoSwap gu2023videoswap preserves the overall look of the source video but fails to swap the reference object (i.e., the green toy car). In contrast, HOI-Swap delivers video edits that realistically depict HOI and faithfully follow the source video’s motion. Beyond this specific example, we note that baseline methods generally perform poorly on our in-contact object swapping task across various examples (see Supp. video for more results). Despite our best efforts to reproduce their results using the official code and adjusting hyperparameters according to the provided guidelines, the performance of these models remained notably suboptimal. We observe similar issues with state-of-the-art text-guided approaches (Supp. C.5). This persistent underperformance highlights a big gap in current (otherwise quite powerful) diffusion models in their ability to accurately model the intricacies of HOIs, indicating great opportunities for further advancements in this field.
@@
Ablation study on motion points sparsity HOI-Swap can generate videos with different degrees of motion alignment with the original video. Figure 5 compares the generated results when sampling no points versus full points. Here the source video depicts a closing trash can lid action. In row 2, when no points are sampled (resulting in no warping), the generated video does not follow the closing lid motion but instead depicts an action of putting the lid into the bin; the visible hand motion outside the inpainting region provides context, guiding the model to generate a plausible sequence with the hand moving downward. Conversely, using all motion points produces videos that closely align with the source video’s motion. See Supp. C.3 for follow-up discussion on this.
Limitations The problem of swapping in-contact objects in videos is inherently very challenging. While HOI-Swap marks an important first step towards addressing these complexities, we recognize its current limitations. Details on these limitations and failure modes are presented in Supp. D.

### v2

@@ 1 Introduction
Consider a video depicting the process of a human hand picking up a kettle and moving it around (Figure 1). What if we want to replace the hand-interacting object in this scene with another item specified by the user—perhaps a differently-shaped kettle, a bottle, or a bowl? This capability—to swap the in-contact object in a video with another, while aligning with the original video’s content—is crucial for enhancing various real-world applications. Such functionality can transform entertainment, allowing users to create novel video content without the need to re-record or engage in labor-intensive manual editing. For example, in advertising, there may be situations where a pre-recorded video needs to adapt to new sponsorship requirements by replacing a soda can in the video with a water bottle. Additionally, it holds significant promise in robotics, where recent results suggest generative models can reduce the reliance on manually collected task-specific visual data and thereby enable large-scale robot learning kapelyukh2023dall ; yang2023learning . For example, imagine a scenario where, from just a single video of a mug being picked up, a generative model is able to produce numerous variants of this video with diverse objects such as bottles, bowls and kettles. This capability could greatly streamline the data collection process, reducing the need for extensive manual data collection.
@@
Second, consider the challenge of spatial alignment with the original video. The reference object might appear in any arbitrary pose; for instance, in Figure 2 (b), the kettle handle is on the left in the reference image, but for realistic interaction, the generated content needs to reposition the kettle handle to the right, where the hand is poised to grasp it. However, current approaches do not offer this level of control, as evidenced by the results from a hand insertion approach Affordance Diffusion (Afford Diff) ye2023affordance in Figure 2 (b). Despite being adept at generating a hand interacting with the given kettle, it does not ensure correct object placement to align with the hand and scene context in the original image, lacking spatial alignment capability.
@@ 3.1 Task Formulation and Framework Overview
We create a fully self-supervised training approach, necessitated by the impracticality of collecting paired videos $(\mathcal{V},\mathcal{V}^{\star})$. We pose the problem as a video inpainting task. Specifically, from the original training video $\mathcal{V}$ and mask $\mathcal{M}$, we derive a masked video $\mathcal{V}^{m}$, accompanied by a set of reference object images $\{\mathcal{I}^{ref}\}$. During training, the model takes the masked video sequence $\mathcal{V}^{m}$ and an image $\mathcal{I}^{ref}$ of the same object originating from a different random timepoint (and hence varying pose and/or viewpoint) in the same training video to reconstruct $\mathcal{V}$. During inference, the model is given a bounding box-masked sequence combined with various object images to test its swapping capability.
As outlined in Section 1, the task presents three main challenges: HOI awareness, spatial and temporal alignment. In response, we propose a two-stage framework that decomposes the complexity. The first stage integrates the reference object into a single frame, targeting HOI awareness and establishing spatial alignment initially. The second stage propagates the one-frame edit across the entire video sequence, focusing on temporal alignment and carrying forward the spatial alignment from the first stage. This structured approach effectively lightens the model’s generation load. In terms of which frame to select from the input video for stage-I edits, we automate the process by adopting an off-the-shelf hand-object detector shan2020understanding to identify frames with hand-object contact, from which we randomly select one for editing. See Supp. C.4 for a detailed analysis. For greater applicability and flexibility, each stage is trained separately in a self-supervised manner, using the original object image as $\mathcal{I}^{ref}$. Consequently, the stage-I model not only plays a crucial role in the video editing pipeline but also serves as a standalone image editing tool. The full pipeline is illustrated in Figure 3.
@@ 3.2 Stage I: HOI-aware Object Swapping in One Frame
  • Object augmentation: Directly forming a training pair ($\mathcal{I}_{i}^{obj}$, $\mathcal{I}_{i}^{m}\rightarrow\mathcal{I}_{i}$) is suboptimal as it does not reflect the variability expected in real-world applications. At test time, the reference object may appear in any pose, orientation, or size. To bridge this gap between training and testing scenarios, we apply strong augmentation techniques: (1) Spatially, we enhance the object’s diversity by applying random resizing, flips, rotations, and perspective transformations; (2) Temporally, instead of using the object image from its original frame $i$, we introduce variability by randomly selecting an alternate frame, $\mathcal{I}_{i^{\prime}}^{obj}$, from all available frames in the source video that contain the object. Finally, for preservation of the reference object’s identity and structural details, we collage the augmented reference object image $\mathcal{I}_{i^{\prime}}^{obj}$ onto the masked region in $\mathcal{I}_{i}^{m}$, producing $\mathcal{I}_{i}^{m^{\prime}}$.
@@ 3.3 Stage II: Controllable Motion-guided Video Generation
In stage II, given the first-stage edit $\mathcal{I}_{i}^{\star}$, source video $\mathcal{V}$, and its binary mask sequence $\mathcal{M}$, the objective is to generate a new video $\mathcal{V}^{\star}$ that propagates the single-frame edit across the remaining frames. For this purpose, we perform warping to transfer pixel points from the edited frame to the remaining frames, resulting in a sparse and incomplete video sequence. We then train a video diffusion network that learns to correct and fill in the gaps, completing the video editing process.
@@
  • Conditioning frame selection. To avoid limiting the model to generating videos based on a specific reference frame, such as the first frame as often used in existing image-to-video generative models, we randomly select a frame $\mathcal{I}_{c}$ from $\mathcal{V}$ as the conditioning frame. This any-frame-conditioning mechanism brings additional flexibility during inference (for detailed discussion, see Supp. C.4).
@@
We vary the sparsity level of sampled motion points, $r$ from 0 to 100 during training, preparing the model for a range of scenarios—from no motion guidance to full motion information. During inference, we demonstrate that the model is capable of generating video sequences with varying levels of motion alignment to the original video (see Section 4.2 and Supp. C.4 for further discussion). Thus, our stage-II training goal is to take the warped yet incomplete video sequence ($\mathcal{V}^{warp}$) as input, then learn to fill in the gaps and smooth over discontinuities to reconstruct the full video ($\mathcal{V}$).
@@
Our video LDM involves two conditional signals in addition to the standard $\mathbf{z}_{t}$ (a noised version of $\mathbf{z}$) and diffusion time step $t$: (1) The warped video sequence $\mathcal{V}^{warp}\in\mathbb{R}^{N\times 3\times H\times W}$, is processed similarly as $\mathcal{V}$, encoded by $\mathcal{E}$, producing a latent feature $\mathbf{z}^{w}\in\mathbb{R}^{N\times 4\times H/8\times W/8}$. $\mathbf{z}^{w}$ is then concatenated channel-wise with $\mathbf{z}$ before being fed into the denoising network $\epsilon_{\theta_{2}}$. (2) The selected frame $\mathcal{I}_{c}$, besides serving as the anchor in the warping process, is also encoded by CLIP radford2021learning , yielding $d^{c}\in\mathbb{R}^{768}$ that provides high-level contextual guidance and is incorporated in $\epsilon_{\theta_{2}}$ via cross-attention. Additionally, please see Supp. C.4 for an ablation of comparing CLIP and DINO as the object encoder.
@@ 4 Experiments
To comprehensively evaluate HOI-Swap, we consider both the image editing (accomplished by the stage-I model) and video editing (facilitated by the entire pipeline) task. We describe the experimental setup in Section 4.1 and present editing results along with an ablation analysis in Section 4.2.
@@ 4.1 Experimental Setup
Datasets Our training leverages two large-scale egocentric datasets, HOI4D liu2022hoi4d and Ego-Exo4D grauman2023ego , which feature abundant HOIs, making them particularly suitable for exploring this problem. In total, the stage-I model is trained on 106.7K frames and the stage-II model uses 26.8K 2-second video clips for training. The evaluation set for image editing includes 1,250 source images, each paired with four reference object images, resulting in a total of 5,000 edited images. The evaluation set for video editing is composed of 25 source videos, also combined with four object images each, yielding 100 unique edited videos. Note that: (1) For image editing evaluation, we include both hand-present scenarios, which challenge the model’s HOI-aware capabilities, and non-hand-present scenarios (20%) to evaluate its general object-swapping capability. (2) For video editing evaluation, in addition to HOI4D videos, we include challenging in-the-wild videos from EPIC-Kitchens damen2018scaling and third-person videos from TCN Pouring sermanet2017time —datasets not used in training—to assess the model’s zero-shot generalization across diverse scenarios. See Supp. B.1 for complete dataset descriptions.
Baselines For image editing, our stage-I model is compared with two strong image inpainting approaches, PBE yang2023paint and AnyDoor chen2023anydoor , along with a hand-insertion approach Afford Diff ye2023affordance in the HOI domain. For video editing, we consider the following baselines: (1) applying the best image editing baseline to each frame of the video; (2) integrating the best image editing baseline with AnyV2V ku2024anyv2v , a recent video editing approach that takes one edited frame from any black-box image editing model as conditional guidance; (3) VideoSwap gu2023videoswap , the state-of-the-art approach in customized object swapping. Besides these object-image-guided approaches, we include a focused comparison with text-guided approaches for a thorough evaluation. See Supp. C.6 for the discussion.
Figure 4: Qualitative results of HOI-Swap. We compare HOI-Swap with image (left) and video editing approaches (right). The reference object image is shown in the upper left corner of the source image. For image editing, HOI-Swap demonstrates the ability to seamlessly swap in-contact objects with HOI awareness, even in cluttered scenes. For video editing, HOI-Swap effectively propagates the one-frame edit across the entire video sequence while accurately following the source video’s motion, achieving the highest overall quality among all methods. We highly encourage readers to check Supp. C.1 and the project page video for more comparisons.
@@ 4.2 Editing Results
Quantitative evaluation Table 1 compares the performance of HOI-Swap on both image and video editing tasks. For image editing, the two inpainting approaches, PBE yang2023paint and AnyDoor chen2023anydoor struggle to generate the hand correctly, as indicated by hand fidelity scores below 90%. Afford Diff ye2023affordance is developed for the HOI context and also benefits from HOI-specific training data in HOI4D. It produces high-fidelity hands but does not address the editing task adequately, resulting in low contact and hand agreement with the source image. Our proposed HOI-Swap excels in both aspects, achieving superior performance in agreement and fidelity scores. In the human evaluation, HOI-Swap is consistently favored over the other three approaches, achieving a preference rate of 72.1%. For video editing, HOI-Swap outperforms all competing methods across all metrics, including general video quality metrics (subject consistency and motion smoothness) and HOI edit metrics (contact agreement, hand agreement and hand fidelity). It significantly surpasses the three baseline approaches, winning the user study with a selection rate of 86.4%. Recognizing the inherent challenges of the video editing task, we included a survey question asking users if “All edits are of very low quality; none successfully swap the object.” This response was selected at an average rate of 10.9%. See Supp. C.3 for details.
@@
For video editing (Figure 4 right), the source video depicts a hand rotating a toy car. The per-frame approach struggles with temporal consistency of the object identity and produces unnatural HOI motions. AnyV2V ku2024anyv2v fails to propagate the first-frame edit through subsequent frames, leading to inconsistencies. VideoSwap gu2023videoswap preserves the overall look of the source video but fails to swap the reference object (i.e., the green toy car). In contrast, HOI-Swap delivers video edits that realistically depict the HOI and faithfully follow the source video’s motion. Beyond this specific example, we note that baseline methods generally perform poorly on our in-contact object swapping task across various examples (see the project page video for more results). Despite our best efforts to reproduce their results using the official code and adjusting hyperparameters according to the provided guidelines, the performance of these models remained notably suboptimal. We observe similar issues with state-of-the-art text-guided approaches (Supp. C.6). This persistent underperformance highlights a big gap in current (otherwise quite powerful) diffusion models in their ability to accurately model the intricacies of HOIs, indicating great opportunities for further advancements in this field.
@@
Ablation study on motion points sparsity HOI-Swap can generate videos with different degrees of motion alignment with the original video. Figure 5 compares the generated results when sampling no points versus full points. Here the source video depicts a closing trash can lid action. In row 2, when no points are sampled (resulting in no warping), the generated video does not follow the closing lid motion but instead depicts an action of putting the lid into the bin; the visible hand motion outside the inpainted region provides context, guiding the model to generate a plausible sequence with the hand moving downward. Conversely, using all motion points produces videos that closely align with the source video’s motion. See Supp. C.4 for follow-up discussion on this.
Figure 6: Qualitative and quantitative comparisons between a one-stage baseline blattmann2023align and our two-stage HOI-Swap. The one-stage model struggles with preserving the new object’s identify and fails to generate accurate interaction patterns, yielding inferior quantitative performance.
One-stage vs. two-stage design Next we verify the efficacy of our two-stage design, where we train a one-stage model variant that takes the reference object image and the masked frame sequence as input and outputs the edited video. Figure 6 shows both qualitative and quantitative comparisons. The one-stage approach does not yield satisfactory results, specifically failing to preserve the reference object’s identity. This limitation arises from the challenge of simultaneously addressing both spatial and temporal alignment, thereby reinforcing the advantage of our two-stage design.
Moreover, we present a comprehensive discussion on various design choices of HOI-Swap in Supp. C.4 and C.5, covering aspects such as object encoder choice, object masking strategy, editing frame selection strategy, and discussion on sampling regions.
5 Limitations and Future work
The problem of swapping in-contact objects in videos poses significant challenges. While HOI-Swap marks an important first step towards addressing these complexities, we recognize its current limitations. Specifically, we identify three areas for improvement: (1) generalization to new objects. HOI-Swap capably swaps new object instances with HOI awareness (e.g., swap an unseen mug with a differently-shaped bowl). More challenging scenarios involve presenting very different unseen reference object images for swapping. For example, the model needs to accurately depict a hand holding scissors, even though scissors were never part of the training dataset. This scenario requires the model to equip “world knowledge” that extends beyond its training data, enabling it to understand and realistically model how hands interact with a broader variety of objects. (2) generalization to long video sequences with complex HOI. The two-stage pipeline of HOI-Swap is motivated by the observation that HOI interactions remain stable throughout a short video clip. For instance, for a picking up mug sequence, replacing the mug with a bowl can be reliably done in one single frame and propagated across the remaining frames. However, for longer sequences, it is possible that an object undergoes multiple distinct hand interactions that change dynamically over time. This complexity necessitates the development of methods that can capture and model these varied HOI interactions across time. (3) controllability. Our proposed controllable motion guidance allows HOI-Swap to freely choose the degree of motion alignment with the source video in editing. One future work direction is to enhance this controllability with spatial support, allowing the model to specify which regions of the source video should be targeted for motion transfer. See Supp. D for more details on HOI-Swap’s limitations and failure modes, including visual examples.